# Anisotropic Ginzburg-Landau model for superconductivity with five-dimensional operators

### v1

@@ II Model
is the non-minimal covariant derivative and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. The usual covariant derivative in turn is given by $D_{\mu}=\partial_{\mu}-ieA_{\mu}$. In the high-energy physics scenario, the vector $\omega^{\mu}$ promotes the violation of Lorentz invariance by setting a privileged direction in space-time and breaking the equivalence between particle and observer transformations. Assuming that the vector is constant, it implies the independence of space-time position, which yields translational invariance and assures the conservation of momentum and energy. Notice that the product $g\omega^{\mu}$ must have dimension of $(\text{mass})^{-1}$.
@@ IV Final Remarks
Finally, it was evident that under certain conditions controlled by the Lorentz-violating contribution, there exists a first-order phase transition between the superconducting state and an insulating state. It is important to highlight here the Lorentz-violating parameter $W$ plays the role of local impurities that could break the homogeneity and isotropy of the system as a whole.

### v2

@@ II Model
is the non-minimal covariant derivative and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. The usual covariant derivative in turn is given by $D_{\mu}=\partial_{\mu}-ieA_{\mu}$. In the high-energy physics scenario, the vector $\omega^{\mu}$ promotes the violation of Lorentz invariance by setting a privileged direction in space-time and breaking the equivalence between particle and observer transformations. Assuming that the vector is constant, it implies the independence of space-time position, which yields translational invariance and assures the conservation of momentum and energy. The interest in this specific model of Lorentz symmetry violation lies in the fact that the covariant derivative in Eq. (1) contains a Pauli-type coupling term, where the complex scalar field directly couples to the electromagnetic field strength. Notice that the product $g\omega^{\mu}$ must have dimension of $(\text{mass})^{-1}$ and it plays the role of a Lorentz violating induced magnetic moment. In $(2+1)$ dimensions, it has been shown that a similar non-minimal coupling term gives rise to fractional spin statistics even in the absence of the Chern-Simons term Nobre:1999mj ; Furtado:2011zz , which is known to exhibit such a characteristic Semenoff:1988jr ; Dunne:1994uy . Fractional statistics play an important role in the interaction between quantum field theory and condensed matter physics Nobre:1999mj ; Wilczek:1982wy .
@@ IV Final Remarks
Finally, it was evident that under certain conditions controlled by the Lorentz-violating contribution, there exists a first-order phase transition between the superconducting state and an insulating state. It is important to highlight here that the Lorentz-violating parameter $W$ plays the role of local impurities that could break the homogeneity and isotropy of the system as a whole. Also, it is worth to mention that the addition of new terms violating Lorentz symmetry, whether in the pure scalar sector Furtado:2020iom , or through different couplings between the scalar sector and the gauge sector or even direct modifications in the pure gauge sector, as proposed by other types of electrodynamics, should lead to results completely different from those presented here. Such additional terms directly impact the equations for the scalar and gauge fields, from which the quantities calculated here are derived. Specifically, we expect modifications both in the critical magnetic field $H_{c2}$ for a type II superconductor and in the behavior of the magnetic field within the superconductor, i.e., related to the London penetration depth. In fact, this will be a subject of investigation in future works.